OpenStudy (anonymous):

Evaluate this integral:

OpenStudy (anonymous):

\[\int\limits_{}^{} x^{3} * (3+x^{2})^{5/2}\]

OpenStudy (anonymous):

So first I set u=x^3 du=3x^2 * dx du * 1/3 = x^2 * dx

OpenStudy (anonymous):

I tried substituting this into the original problem but it didn't work...

ganeshie8 (ganeshie8):

try \(u = 3+x^2\) instead

OpenStudy (anonymous):

Okay so you would get 3+x^2 = u 2xdx = du xdx = du/2 Can you raise both sides of the equation by the third power? Wouldn't it be (xdx)^3 though? dx^3 is not possible?

ganeshie8 (ganeshie8):

``` 3+x^2 = u 2xdx = du xdx = du/2 ``` so far so good

ganeshie8 (ganeshie8):

next notice that you can write x^3 as x*x^2

ganeshie8 (ganeshie8):

\(3 + x^2 = u\implies x^2 = u-3\) so the integral becomes \[\begin{align}\int\limits_{}^{} x^{3} * (3+x^{2})^{5/2}\,dx &= \int\limits_{}^{} x*x^{2} * (3+x^{2})^{5/2}\,dx \\~\\ &=\int (u-3)(u)^{5/2} \frac{du}{2}\end{align}\]

OpenStudy (anonymous):

ooohoohhhh! I got the answer now! I just didn't see how we could separate x^3 into x*x^2hahaha #simplealgebra... Thank you~

ganeshie8 (ganeshie8):

yw :)